Here is the current status of ITS services after the failure of the HSB core switch today:
- All core service are up and functioning.
- About 95% of the wireless network is functional. The remaining portion will not be functional until we receive the replacement part.
- Some back-office systems located only in BRK are down until the replacement part arrives. We do not know of any end-user services (other than wireless) that are still affected.
- When the new part arrives tomorrow morning, we should be able to install it and bring up the BRK switch with no disruption to the campus services.
- We will need to schedule a maintenance window (30 minutes should be enough) within the next week to restore some network cabling connections to their original locations.
- One of the four units that make up the MPC SAN volumes has apparently failed (Tom is still working on it). This unit hosts a portion of the eDocs and Blackboard data. The data is currently accessible, but we cannot sustain a second failure without loss of all data stored on the devices. We are currently in the process of obtaining a replacement SAN. Until then, Roger will be monitoring the daily backups to ensure that we are protected.
- It took about 6 hours from the time of failure to restore Blackboard and about 7.5 hours to restore all other failed services.